# Structure growth in $f(Q)$ cosmology

### v1

@@ Abstract
We take into account redshift-space distortion measurements to investigate the growth of cosmological large-scale structures within the framework of modified symmetric teleparallel $f(Q)$ gravity. After comparing the predictions of the $f(Q)$-gravity expansion history with OHD and SNIa datasets and constraining the pertinent cosmological parameters $\Omega_{m}$ and $H_{0}$, together with the exponent $n$ for $f(Q)$ power-law models, we derive the full system of equations governing linear cosmological perturbations to study matter fluctuations using the $1+3$ covariant formalism when applied to $f(Q)$ gravity. Thus, we resort to both the usual redshift-space distortion data $f\sigma_{8}$ and some recent separate measurements of the growth rate $f$ and the amplitude of matter fluctuations $\sigma_{8}$ from the VIPERS and SDSS collaborations to find the best-fit cosmological parameters $\Omega_{m}$, $\sigma_{8}$ and $n$. We also apply a collective analysis of such growth-structure data together with the aforementioned cosmic expansion measurements, to restrict these parameters through Monte Carlo Markov Chain simulations. As a result, our analysis is capable of determining the statistical significance for the best-fit parameter values through the AIC and BIC Bayesian selection criteria.
@@ 1 Introduction
Extensions of general relativity (GR) have drawn significant interest in explaining the late-time accelerated expansion history of the universe [1] with no requirement for a scenario that includes dark energy ad hoc. Such attempts have also drawn great attention in the realm of relativistic cosmology to solve many unsolved problems such as inhomogeneities and anisotropies, matter density fluctuations, and the formation of large-scale structures, to mention just a few. The recent observational data presented in [2] showed that our observable universe is experiencing an accelerating expansion, perhaps the clearest clue to the properties of gravity beyond GR. The main pillar of the mystery of cosmic expansion is that a major part of the universe is made up of a substance known as dark energy [3] or extended - also dubbed modified - gravity on cosmological scales. The Einstein-Hilbert action is described by $f(Q)$ $Q$ being the nonmetricity scalar [4], $f(T)$ $T$ being the torsion scalar [5], $f(R)$ $R$ being the Ricci scalar [6], a combination of the two dubbed $f(R,T)$ [7] or $f(G)$ $G$ being the Gauss-Bonnet invariant [8], among others.
@@
One of the recent additions to the myriad of modifications to GR, $f(Q)$-gravity theory, is based on the generalization of the symmetric teleparallel equivalent of GR (STEGR), characterized by a connection in which both the curvature and torsion vanish. Consequently, STEGR is described by a nonmetric scalar $Q$ which represents one of the equivalent geometric formulations of GR [9]. Thus, $f(Q)$ has gained much attention to study different aspects of cosmology and parameter constraints [10]. The detailed clarifications of STEGR when referring to the background evolution of the universe have been studied in [11] whereas the different self-accelerating solutions have been studied in [12], all of resorting to the predictions of the cosmological expansion history within this class of theories resorting solely to observations such as high-redshift Hubble diagrams from SNIa, baryon acoustic oscillations (BAO) or CMB shift factor, all of them based on different distance measurements which are only sensitive to the expansion history.
@@
Nevertheless, there exist a different type of observations that are sensitive to the expansion history but also to the evolution of matter density perturbations. The fact that the evolution of perturbations depends on the specific gravity theory, i.e., it differs in general from that of Einstein’s gravity even though the background evolution is the same, has made this kind of observation instrumental in distinguishing between different models and mechanisms causing the cosmological acceleration. Within this context, there exist two basic approaches to cosmological perturbations in the literature: the first is the gauge-invariant perturbations, also dubbed metric, formalism [13], whereas the second approach is the $1+3$ covariant gauge-invariant perturbations formalism [14] which will be employed in this work. The first approach presents several shortcomings such that it does not provide a covariant description [15], and can be nonlinear because it is built around a local comoving observer frame. The second approach though is covariant and gauge invariant. Also, as widely known, it uses the kinematic quantities, the energy-momentum tensor of the fluid(s), and the gravito-electromagnetic parts of the Weyl tensor Also, the second approach to perturbation theory produces differential equations that help to describe the true physical degrees of freedom. As an example, different works have used the $1+3$ covariant formalism in GR [16], $f(R)$ [17, 18] and, in the context of $f(T)$ to investigate the evolution of large-scale structures [19, 20] and to compute the gravitational instabilities seeding the formation of stars, quasars, galaxies and clusters [21].
@@
Amongst those cosmological observations that are sensitive to the evolution of matter density perturbations, the fact that the spatial distribution of galaxies appears squashed and distorted [22] when their positions are shown as functions of their redshift rather than as functions of their distance, plays nowadays an instrumental role in observational cosmology. This is the well-known redshift-space distortion effect. It is possible to measure redshift space distortions to set constraints on cosmological parameters, and as a byproduct, it is also possible to integrate these distortions to recoup the underlying actual space correlation function. In this regard, data on growth rate measurements $f\sigma_{8}=\sigma_{8}\frac{\rm d\ln{\mathcal{D}}}{\rm d\ln{a}}$ have recently been disclosed, where $a$ is the cosmological scale factor, $\mathcal{D}$ is the linear growth of matter overdensity and $\sigma_{8}$ is the normalization of the matter power spectrum on the scale of $8h^{-1}{\rm Mpc}$. As a consequence, such growth-structure data have been considered to study different modified theories of gravity. For instance, just to cite a few efforts, in [23] the growth of structures in $f(Q)$ gravity has been performed under the metric formalism by following the designer approach. In [24] $N$-body simulations in $f(R)$ gravity enabled the authors to make redshift-space predictions on the dark-matter clustering such theories. Finally, the use of growth structure data in the context of the effective field theory of dark energy formalism presented allowed authors in [25] to constrain the cosmological parameters.
@@
In this work, we shall first pay attention to the cosmological expansion history resorting to both OHD, SNIa and combined OHD+SNIa data to constrain the cosmological parameters for power-law gravity $f(Q)=Q+F(Q)$ models [9], where $F(Q)=\alpha Q^{n}$ [^1] For the case of $n=0$, this gravity model reduces to $\Lambda$CDM.. In this manuscript, both the MCMC hammer [26] and GetDist [27] software packages shall be used to constrain these best-fit parameters $\Omega_{m}$, $H_{0}$, $n$} for $f(Q)$ gravity model by using the SNIa distance modulus measurements from the Pantheon analysis [28], which consists of 1048 distinct SNIa ranging in the redshift interval $0.01012\leq z\leq 2.26$. We shall also use the observational Hubble parameter data (OHD), from Baryon acoustic oscillations (BAO) provided by the Sloan Digital Sky Survey (SDSS), DR9, and DR11 and cosmic chronometers or differential ages of galaxies at the redshifts range $0.0708<z\leq 2.36$, [29].
@@
Second, after studying the background cosmology in the framework of $f(Q)$ gravity, the crucial point in this investigation is to perform a detailed analysis of linear cosmological perturbations using the $1+3$ covariant formalism to study the growth of large-scale structures. Therefore we shall consider both the whole system and the so-called quasistatic approximation of the perturbation equations. This will allow us to better understand the limitations of the quasi-static approximation when studying the growth density fluctuations in the realm of $f(Q)$ theories. For further tasks, we have implemented the sets of redshift-space distortions data $f\sigma{{}_{8}}$ with the three latest separate measurements of the growth rate $f$ and the amplitude of matter fluctuations $\sigma_{8}$ from the VIPERS and SDSS collaborations. In particular, we will compare constraints obtained from three data sets: $i)$ the combinations of the measurements of the growth rate ($f$), and $\sigma_{8}$, sourced from SDSS and Vipers PDR-2 Collaborations [30], collectively labeled as $\sigma_{8}+f$; $ii)$ the 30 redshift-distortion measurements of $f\sigma_{8}$, dubbed $f\sigma_{8}$, covering the redshift range $0.001\leq z\leq 1.944$ and; $iii)$ the full combinations of thirty $f\sigma+8$, three combined three $f$, and three $\sigma_{8}$ measurements collectively labeled as $\sigma_{8}+f+f\sigma_{8}$ (for further details refer to Table 6 and 6 in Appendix A). For the latter case $iii)$, the $f\sigma+8$ measurements with their $f$ and $\sigma_{8}$ counterparts are indeed not taken into account for the independence of the data. Using MCMC simulations, these sets of data - either by themselves or combined with OHD and SNIa measurements - allowed us to constrain the corresponding best fit values of $\sigma_{8}$, $\Omega_{m}$, and $n$ for $f(Q)$ gravity power-law models.
@@
The layout of the manuscript is as follows: in Section 2, the cosmology of $f(Q)$ gravity is reviewed; the cosmological parameters are constrained and the viability of the $f(Q)$ gravity model is statistically analyzed using OHD, SNIa and OHD+SNIa datasets. Then, in Section 3 the $f(Q)$-gravity first- and second-order linear evolution equations using the $1+3$ covariant formalism are derived. This section applies scalar and harmonic decomposition techniques to find solutions for density fluctuations. We also present the theoretical results of the matter density fluctuations when applied to the power-law $f(Q)$-gravity models. Subsequently, Section 4 is devoted to comparing the linear growth of the matter density predictions against $\sigma_{8}+f$, $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$ and presents the statistical significance for $f(Q)$-gravity power law models. Also, for these models, an exhaustive joint analysis of theoretical predictions against growth structure data and cosmic expansion measurements, i.e., OHD and SNIa data, is performed here. Finally, we conclude with the conclusions of the manuscript in Section 5.
@@ 2 Background cosmology of $f(Q)$-gravity
The action of non-metric gravity as presented in [9]
$S=\int\sqrt{-g}\left(\frac{1}{2}f(Q)+\mathcal{L}_{m}\right){\rm d}^{4}x,$ (2.1)
where $f(Q)$ is an arbitrary function of the nonmetricity $Q$, ${g}$ is the determinant of the metric $g_{\mu\nu}$ and $\mathcal{L}_{m}$ is the matter Lagrangian density [^2] Unless otherwise indicated, throughout the manuscript we use $c=8\pi G=1$ units.. Variations of action (2.1) with respect to the metric tensor yields the field equations as
$\frac{1}{\sqrt{-g}}\nabla_{\gamma}\left(\sqrt{-g}f_{Q}P^{\gamma}_{\mu\nu}\right)+\frac{1}{2}g_{\mu\nu}f+f_{Q}\left(P_{\mu\nu\delta}Q_{v}^{\gamma\delta}-2Q_{\gamma\delta\mu}P^{\gamma\delta}_{\nu}\right)=-T_{\mu\nu}\;,$
@@
$Q_{\gamma\mu\nu}\equiv\nabla_{\gamma}g_{\mu\nu}\;,$ (2.3)
@@
$Q_{\gamma}=Q_{\gamma}~{}^{\mu}~{}_{\mu}\qquad\tilde{Q}_{\gamma}=Q^{\mu}~{}_{\gamma\mu}\,,$ (2.4)
@@
$P^{\gamma}_{\mu\nu}=\frac{1}{4}\left(-Q^{\gamma}~{}_{\mu\nu}+2Q_{(\mu}~{}^{\gamma}~{}_{\nu)}-Q^{\gamma}g_{\mu\nu}-\tilde{Q}^{\gamma}g_{\mu\nu}-\delta^{\gamma}_{(\mu}Q_{\nu)}\right)\;,$ (2.5)
@@
$\qquad Q=-Q_{\mu\nu\gamma}P^{\mu\nu\gamma}\;.$ (2.6)
@@
${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t)\delta_{\alpha\beta}{\rm d}x^{\alpha}{\rm d}x^{\beta}\;,$ (2.7)
[…]
$3H^{2}=\rho_{m}+\rho_{r}+\frac{F}{2}-QF^{\prime}\;,$ (2.8)
$2\dot{H}+3H^{2}=-\frac{\rho_{r}}{3}-\frac{F}{2}+QF^{\prime}+\frac{H\dot{Q}}{Q}(2QF^{\prime\prime}+F^{\prime})\;,$ (2.9)
respectively, where $\rho_{m}$ and $\rho_{r}$, the energy density of non-relativistic matter and (relativistic matter (radiation)  [^4] Since the contribution of radiation to the late-time cosmological expansion history is so minimal, we have safely neglected such a contribution in the analysis to perform., $F^{\prime}$ and $F^{\prime\prime}$ are the first and second derivative of $F$ with respect to $Q$ respectively, and the thermodynamic quantities for the effective $f(Q)$ fluid are defined as
$\rho_{Q}=\frac{F}{2}-QF^{\prime},$ (2.10)
$p_{Q}=-\rho_{Q}+\frac{\theta\dot{Q}}{3Q}(2QF^{\prime\prime}+F^{\prime})\;,$ (2.11)
where $\theta=3H$. In the rest of this investigation, we use a generic power-law $F(Q)$-gravity model
$F(Q)=\alpha\,Q^{n}\;,$ (2.12)
where $n$ is a constant, for $n=0$, the function $f(Q)$ reduces to $\Lambda$CDM. Similarly, when $n=1/2$, the above modified Friedmann and Raychaudhuri equations are simplified to GR. From Eq. (2.8), we obtain $\alpha$ [^5] In the following we shall assume $n\neq 1/2$, i.e., $f(Q)$ power-law models strictly different from pure GR ($\Omega_{m}=1$) in order to avoid expression (2.13) being indeterminate. as
$\alpha=\frac{1-\Omega_{m}}{(6H_{0}^{2})^{n-1}(1-2n)}\;.$ (2.13)
Then with the choice in (2.12) and using (2.13), the normalized Hubble parameter - for $n\neq 1/2$ yields,
$E^{2}\equiv\frac{H^{2}(z)}{H_{0}^{2}}={\Omega_{m}}(1+z)^{3}+(1-\Omega_{m})E^{2n}\,,$ (2.14)

### v4

@@ Abstract
We take into account redshift space distortion measurements to investigate the growth of cosmological large-scale structures within the framework of generalised symmetric teleparallel - $f(Q)$ gravity. After comparing the predictions of the $f(Q)$-gravity expansion history with OHD and SNIa Pantheon+ sample datasets and constraining the pertinent cosmological parameters $\Omega_{m}$ and $H_{0}$, together with the exponent $n$ for $f(Q)$ power-law models, we derive the full system of equations governing linear cosmological perturbations to study matter fluctuations using the $1+3$ covariant formalism when applied to $f(Q)$ gravity. Thus, we resort to both the usual redshift-space distortion data $f\sigma_{8}$ and some recent separate measurements of the growth rate $f$ and the amplitude of matter fluctuations $\sigma_{8}$ from the VIPERS and SDSS collaborations to find the best-fit cosmological parameters $\Omega_{m}$, $\sigma_{8}$ and $n$. We also apply a collective analysis of such growth-structure data together with the aforementioned cosmic expansion measurements, to constrain these parameters through Monte Carlo Markov Chain simulations. As a result, our analysis is capable of determining the statistical significance for the best-fit parameter values through the AIC and BIC Bayesian selection criteria.
keywords:
cosmological parameters - dark energy – cosmology: observations – cosmology: theory.
pubyear: 2025pagerange: Structure growth in $f(Q)$ cosmology0.00LABEL:lastpage
@@ 1 Introduction
Extensions of general relativity (GR) have drawn significant interest in explaining the late-time accelerated expansion history of the universe (Riess et al., 1998) with no requirement for a scenario that includes dark energy ad hoc. Such attempts have also drawn great attention in the realm of relativistic cosmology to solve many unsolved problems such as inhomogeneities and anisotropies, matter density fluctuations, and the formation of large-scale structures, to mention just a few. The recent observational data presented in (Reichardt et al., 2012) showed that our observable universe is experiencing an accelerating expansion, perhaps the clearest clue to the properties of gravity beyond GR. The main pillar of the mystery of cosmic expansion is that a major part of the universe is made up of a substance known as dark energy (Barris et al., 2004) or extended - also dubbed modified - gravity on cosmological scales. The Einstein-Hilbert action is described by $f(Q)$ $Q$ being the nonmetricity scalar (Capozziello & D’Agostino, 2022), $f(T)$ $T$ being the torsion scalar (Cai et al., 2016), $f(R)$ $R$ being the Ricci scalar (De la Cruz-Dombriz & Dobado, 2006), a combination of the two dubbed $f(R,T)$ (Myrzakulov, 2012) or $f(G)$ $G$ being the Gauss-Bonnet invariant (De Felice & Tsujikawa, 2009), among others.
@@
One of the recent additions to the myriad of modifications to GR, $f(Q)$-gravity theory is based on the generalisation of the symmetric teleparallel equivalent of GR (STEGR), characterised by a connection in which both the curvature and torsion vanish. Consequently, STEGR is described by a non-metric scalar $Q$, representing one of the equivalent geometric formulations of GR (Khyllep et al., 2021). Thus, $f(Q)$ has gained much attention to study different aspects of cosmology and parameter constraints (D’Agostino & Nunes, 2022). The detailed clarifications of STEGR when referring to the background evolution of the universe have been studied in (Ayuso et al., 2021) whereas the different self-accelerating solutions have been studied in (Junior & Rodrigues, 2023), all of resorting to the predictions of the cosmological expansion history within this class of theories resorting solely to observations such as high-redshift Hubble diagrams from SNIa, baryon acoustic oscillations (BAO) or CMB shift factor, all of them based on different distance measurements which are only sensitive to the expansion history.
@@
Nevertheless, there exist a different type of observations that are sensitive not only to the expansion history but also to the evolution of matter density perturbations. The fact that the evolution of perturbations depends on the specific gravity theory, i.e., it differs in general from that of Einstein’s gravity even though the background evolution is the same, has made this kind of observations instrumental in distinguishing between different cosmological models and the mechanisms that cause cosmological acceleration. Within this context, there exist two basic approaches to cosmological perturbations in the literature: the first is the gauge-invariant perturbations, also dubbed metric, formalism (Kodama & Sasaki, 1984), whereas the second approach is the $1+3$ covariant gauge-invariant perturbations formalism (Dunsby et al., 1992) which will be employed in this work. The first approach presents several shortcomings such that it does not provide a covariant description (Tsagas et al., 2008), and can be nonlinear because it is built around a local comoving observer frame. The second approach though is covariant and gauge invariant. Also, as widely known, it uses the kinematic quantities, the energy-momentum tensor of the fluid(s), and the gravito-electromagnetic parts of the Weyl tensor. Also, the second approach to perturbations theory produces differential equations that help to describe the true physical degrees of freedom. As an example, different works have used the $1+3$ covariant formalism in GR (Dunsby, 1991), $f(R)$ (Abebe et al., 2012; Ntahompagaze et al., 2018) and, in the context of $f(T)$ (Sami et al., 2021; Sahlu et al., 2020) to investigate the evolution of large-scale structures and to compute the gravitational instabilities seeding the formation of stars, quasars, galaxies, and clusters (Fry, 1984).
@@
Amongst those cosmological observations that are sensitive to the evolution of matter density perturbations, the fact that the spatial distribution of galaxies appears squashed and distorted (Percival et al., 2011) when their positions are shown as functions of their redshift rather than as functions of their distance, plays an instrumental role in observational cosmology nowadays. This is the well-known redshift-space distortion effect. It is possible to measure redshift space distortions to set constraints on cosmological parameters, and as a byproduct, it is also possible to integrate these distortions to recoup the underlying actual space correlation function. In this regard, data on growth rate measurements $f\sigma_{8}=\sigma_{8}\frac{\rm d\ln{\mathcal{D}}}{\rm d\ln{a}}$ have recently been disclosed, where $a$ is the cosmological scale factor, $\mathcal{D}$ is the linear growth of matter overdensity and $\sigma_{8}$ is the normalisation of the matter power spectrum on the scale of $8h^{-1}{\rm Mpc}$. As a consequence, such growth-structure data have been considered to study different modified theories of gravity. For instance, just to cite a few efforts, in (Albuquerque & Frusciante, 2022) the growth of structures in $f(Q)$ gravity has been performed under the metric formalism by following the designer approach. In (Jennings et al., 2012) $N$-body simulations in $f(R)$ gravity enabled the authors to make redshift-space predictions on the dark-matter clustering such theories. Finally, the use of growth structure data in the context of the effective field theory of dark energy formalism presented allowed authors in (Perenon et al., 2019) to constrain the cosmological parameters.
@@
In this work, we shall first pay attention to the cosmological expansion history resorting to both OHD, SNIa and combined OHD+SNIa data to constrain the cosmological parameters for power-law gravity $f(Q)=Q+F(Q)$ models (Khyllep et al., 2021), where $F(Q)=\alpha Q^{n}$. For the case of $n=0$, this gravity model reduces to $\Lambda$CDM. In this communication, both the MCMC hammer (Foreman-Mackey et al., 2013) and GetDist (Lewis, 2019) software packages shall be used to constrain these best-fit parameters $\Omega_{m}$, $H_{0}$, $n$} for $f(Q)$ gravity model by using SNIa distance moduli measurements from the Pantheon+ sample Brout et al. (2022), which consists of 1701 light curves of 1550 distinct SNIa ranging in the redshift interval $z\in[0.001,2.26]$. We shall also use the observational Hubble parameter data (OHD), from Baryon acoustic oscillations (BAO) provided by the Dark Energy Spectroscopic Instrument (DESI) 2024 Survey Adame et al. (2024) and cosmic chronometers or differential ages of galaxies at the redshift range $0.0708<z\leq 2.36$, (Yu et al., 2018).
@@
Second, after studying the background cosmology in the framework of $f(Q)$ gravity, the crucial point in this investigation is to perform a detailed analysis of linear cosmological perturbations using the $1+3$ covariant formalism to study the growth of large-scale structures. Therefore, we shall consider both the whole system and the so-called quasistatic approximation of the perturbation equations. This will allow us to better understand the limitations of the quasi-static approximation when studying the growth density fluctuations in the realm of $f(Q)$ theories. For further tasks, we have implemented the sets of redshift-space distortion data $f\sigma{{}_{8}}$ with the three latest separate measurements of the growth rate $f$ and the amplitude of matter fluctuations $\sigma_{8}$ from the VIMOS Public Extragalactic Redshift Survey (VIPERS) and SDSS collaborations. In particular, we will compare constraints obtained from three datasets: $i)$ the combinations of the measurements of the growth rate ($f$), and $\sigma_{8}$, sourced from SDSS and Vipers PDR-2 Collaborations (De La Torre et al., 2017), collectively labeled as $\sigma_{8}+f$; $ii)$ the 30 redshift-distortion measurements of $f\sigma_{8}$, dubbed $f\sigma_{8}$, covering the redshift range $0.001\leq z\leq 1.944$ and; $iii)$ the full combinations of thirty $f\sigma_{8}$, three combined three $f$, and three $\sigma_{8}$ measurements collectively labeled as $\sigma_{8}+f+f\sigma_{8}$ (for further details refer to Table 6 and 6 in Appendix A). For the latter case $iii)$, the $f\sigma+8$ measurements with their $f$ and $\sigma_{8}$ counterparts are indeed not taken into account for the independence of the data. Using MCMC simulations, these sets of data - either by themselves or combined with OHD and SNIa measurements - will allow us to constrain the corresponding best-fit values of $\{H_{0},\Omega_{m},\sigma_{8},n\}$. In order to do so, throughout this manuscript, the following priors have been assumed: $H_{0}=[50.00,80.00]$ in ${\rm km\,s^{-1}\,Mpc^{-1}},\,\Omega_{m}=[0.00,1.00]$, $\sigma_{8}=[0.50,1.00]$ and $n=[-0.50,0.50]$, for the power-law $f(Q)=Q+\alpha Q^{n}$ gravity model under study in what follows.
At this stage, a pertinent comment refers to the crucial difficulty teleparallel gravity theories, namely $f(G)$ and $f(T)$ and $f(Q)$, face in balancing strong couplings due to ghost instabilities. For instance, in the context of $f(Q)$ theories, the choice of connection affects the number of degrees of freedom and the overall stability. Such a choice can help reduce the likelihood of ghost formation. As discussed in Aguiar Gomes et al. (2024), two of those connections [^1] In this communication we have chosen one of those two, $C_{1}$ following the terminology in Aguiar Gomes et al. (2024). exhibit suppressed linear spectra, indicating they are strongly coupled. For the remaining connection, such an analysis disclosed the presence of seven gravitational degrees of freedom, with at least one identified as a ghost. A potential solution to these problems could involve exploring a direct coupling between the connection and matter fields, or a non-minimal coupling, as suggested by Heisenberg et al. (2024) and exemplified in Bello-Morales et al. (2024). In our study, we have decided nonetheless to consider a gravitational action with minimal coupling between the matter and the connection, as described in Eq. (1) below, even if ghost issues may emerge. Our goal is thus twofold: first to provide a full derivation of the $1+3$ formalism applicable to scalar perturbations for $Q$-based theories, which can be easily extended for different couplings; and second to conclude, that even without invoking the ghost’s presence, the competitiveness of power-law $f(Q)$ models against $\Lambda$CDM remains limited.
@@
The layout of the manuscript is as follows: in Section 2, the cosmology of $f(Q)$ gravity is reviewed; the cosmological parameters are constrained, and the viability of the $f(Q)$ gravity model is statistically analysed using OHD, SNIa, and OHD+SNIa datasets. Also, some caveats of these models are presented. Then, in Section 3 the $f(Q)$-gravity first- and second-order linear evolution equations using the $1+3$ covariant formalism are derived. This section applies scalar and harmonic decomposition techniques to find solutions for density fluctuations. We also present the theoretical results of the matter density fluctuations when applied to the power-law $f(Q)$-gravity models. Subsequently, Section 4 is devoted to comparing the linear growth of the matter density predictions against $\sigma_{8}+f$, $f\sigma_{8}$ and $\sigma_{8}+f+f\sigma_{8}$ and presents the statistical significance for $f(Q)$-gravity power-law models. Also for these models, an exhaustive joint analysis of theoretical predictions against growth structure data and cosmic expansion measurements, i.e., OHD and SNIa data, is performed here. Finally, we conclude with the conclusions of the manuscript in Section 5.
@@ 2 Background cosmology of $f(Q)$-gravity
The action of non-metric gravity as presented in (Khyllep et al., 2021; Capozziello et al., 2022)
$S=\int\sqrt{-g}\left(\frac{1}{2}f(Q)+\mathcal{L}_{m}\right){\rm d}^{4}x,$ (1)
where $f(Q)$ is an arbitrary function of the nonmetricity $Q$, ${g}$ is the determinant of the metric $g_{\mu\nu}$ and $\mathcal{L}_{m}$ is the matter Lagrangian density [^2] Unless otherwise indicated, throughout the manuscript we use $c=8\pi G=1$ units.. Variations of action (1) with respect to the metric tensor yields the field equations as
$\frac{{2}}{\sqrt{-g}}\nabla_{\gamma}\left(\sqrt{-g}f_{Q}P^{\gamma}_{\mu\nu}\right)+\frac{1}{2}g_{\mu\nu}f$
$+f_{Q}\left(P_{\mu\nu\delta}Q_{v}^{\gamma\delta}-2Q_{\gamma\delta\mu}P^{\gamma\delta}_{\nu}\right)=-T_{\mu\nu}\;,$
@@
$Q_{\gamma\mu\nu}\equiv\nabla_{\gamma}g_{\mu\nu}\;,$ (3)
@@
$Q_{\gamma}=Q_{\gamma}~{}^{\mu}~{}_{\mu}\qquad\tilde{Q}_{\gamma}=Q^{\mu}~{}_{\gamma\mu}\,,$ (4)
@@
$P^{\gamma}_{\mu\nu}=\frac{1}{4}\Bigg{(}-Q^{\gamma}~{}_{\mu\nu}+2Q_{(\mu}~{}^{\gamma}~{}_{\nu)}-Q^{\gamma}g_{\mu\nu}$
$-\tilde{Q}^{\gamma}g_{\mu\nu}-\delta^{\gamma}_{(\mu}Q_{\nu)}\Bigg{)}\;,$ (5)
@@
$\qquad Q=-Q_{\mu\nu\gamma}P^{\mu\nu\gamma}\;.$ (6)
@@
${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t)\delta_{\alpha\beta}{\rm d}x^{\alpha}{\rm d}x^{\beta}\;,$ (7)
where $a(t)$ is the cosmological scale factor and $\delta_{\alpha\beta}$ the Kronecker delta. Thus, the trace of the nonmetric tensor becomes $Q=6H^{2}$ and the corresponding modified Friedmann and Raychaudhuri equations, once the gravitational Lagrangian has been split as $f(Q)=Q+F(Q)$ become
$3H^{2}=\rho_{m}+\rho_{r}+\frac{F}{2}-QF^{\prime}\;,$ (8)
$2\dot{H}+3H^{2}=-\frac{\rho_{r}}{3}-\frac{F}{2}+QF^{\prime}+\frac{H\dot{Q}}{Q}(2QF^{\prime\prime}+F^{\prime})\;,$ (9)
respectively. Here $\rho_{m}$ and $\rho_{r}$ are the energy density of non-relativistic matter and (relativistic matter (radiation)  [^3] Since the contribution of radiation to the late-time cosmological expansion history is so minimal, we have safely neglected such a contribution in the analysis to perform., $F^{\prime}$ and $F^{\prime\prime}$ are the first and second derivative of $F$ with respect to $Q$ respectively, and the thermodynamic quantities for the effective $f(Q)$ fluid are defined as
$\rho_{Q}=\frac{F}{2}-QF^{\prime},$ (10)
$p_{Q}=-\rho_{Q}+\frac{\theta\dot{Q}}{3Q}(2QF^{\prime\prime}+F^{\prime})\;,$ (11)
where $\theta\equiv 3H$. In the rest of this investigation, we use a generic power-law $F(Q)$-gravity model
$F(Q)=\alpha\,Q^{n}\;,$ (12)
where $n$ is a constant such that for $n=0$, the function $f(Q)$ reduces to $\Lambda$CDM. Similarly, when $n=1/2$, the above modified Friedmann and Raychaudhuri equations are simplified to those of GR. From Eq. (8), we obtain $\alpha$ [^4] In the following we shall assume $n\neq 1/2$, i.e., $f(Q)$ power-law models strictly different from matter-only GR ($\Omega_{m}=1$) in order to avoid expression (13) being indeterminate. as
$\alpha=\frac{1-\Omega_{m}}{(6H_{0}^{2})^{n-1}(1-2n)}\;.$ (13)
Then with the choice in (12) and using (13), the normalized Hubble parameter - for $n\neq 1/2$ yields,
$E^{2}\equiv\frac{H^{2}(z)}{H_{0}^{2}}={\Omega_{m}}(1+z)^{3}+(1-\Omega_{m})E^{2n}\,,$ (14)